for the function y=x^2 - 4x -5 how do you put it in the form y+a(x-h)^2 + k
Answer by mukhopadhyay(490)

y=x^2 - 4x -5
=> y = x^2 - 4x + 4 - 9
=> y = (x-2)^2 - 9
=> y - (x-2)^2 + 9 = 0
The above is in the same form as described where a = -1, h=2, and k=9